# Decentralized Cooperation in Heterogeneous Multi-Agent Reinforcement Learning via Graph Neural Network-Based Intrinsic Motivation

### v1

@@ Introduction
Table 3: Robustness of CoHetteam against increasing number of agents in VMAS-Navigation
Conclusions and Future Work
In this paper, we propose the CoHet algorithm, a novel intrinsic reward calculation mechanism that utilizes a Graph Neural Network (GNN) and is suitable for decentralized heterogeneous MARL policy learning algorithms. It can be employed as a standalone module alongside existing decentralized policy optimization methods. Our proposed reward calculation architecture is applicable to practical multi-agent systems that operate under the constraints of partial observability and reward sparsity. The algorithm is validated in a series of cooperative multi-agent tasks that involve heterogeneous agents. We augment our research with an exploration of the effects of the agent dynamics model on the intrinsic reward module, compare the two variants of the CoHet algorithm and demonstrate that CoHetteam is robust against an increase in heterogeneous agents in the environment.
A promising direction for future work includes exploring the alternative types of intrinsic motivation such as curiosity-driven or novelty-based rewards in the context of decentralized heterogeneous learning frameworks. Finding the proper balance between intrinsic and extrinsic rewards remains an open research problem. Future works can explore utilizing different weighting mechanisms that prioritize the predictions of the agents with the same sub-goals, and heterogeneity types. Ultimately, we opine that the collaboration between the agents in a MARL setting should take the need for decentralized training and agent heterogeneity into account.

### v2

@@ 5.4. Reward Architecture Evaluation
5.5. Robustness to Increasing No. of Agents
6. Conclusions and Future Work
In this paper, we introduce CoHet, a novel intrinsic reward mechanism leveraging a Graph Neural Network (GNN) for decentralized heterogeneous MARL policy learning. CoHet can be integrated with existing decentralized policy optimization methods and is well-suited for practical multi-agent systems facing partial observability and reward sparsity. We validate CoHet across cooperative multi-agent tasks involving heterogeneous agents, explore the impact of agent dynamics on the intrinsic reward module, and show that CoHetteam remains robust as the number of heterogeneous agents increases.
A promising direction for future work includes exploring the alternative types of intrinsic motivation such as curiosity-driven or novelty-based rewards in the context of decentralized heterogeneous policy learning frameworks. Finding the proper balance between intrinsic and extrinsic rewards remains an open research problem. Future works can explore utilizing other types of weighting mechanisms that prioritize the predictions of the agents with the same sub-goals, and heterogeneity types. Ultimately, we opine that the collaboration between the agents in a MARL setting should take the need for decentralized training and agent heterogeneity into account.